OpenStudy (anonymous):

The distribution of heights of a population of adults is approximately normal with mean 66 inches and SD 2.5 inches. Approximately what percent of the adults have heights that are within 1 inch of the average

OpenStudy (anonymous):

please guide me

OpenStudy (nurali):

1/2.5 = 0.4. Enter z=-0.4 into the lower endpoint and z=0.4 into the upper endpoint on the SticiGui applet to find 31.08%.

OpenStudy (nurali):

if correct please click best response.

OpenStudy (anonymous):

approximately what percent of the adults are 70 inches tall, to the nearest inch?now tell this

OpenStudy (nurali):

Adults that are 70 inches tall to the nearest inch have heights ranging from 69.5 to 70.5. In standard units,(69.5 - 66)/2.5 = 1.4; (70.5 - 66)2.5 = 1.8. Enter these z-values into the lower and upper endpoints respectively to find that the area is equal to 4.48%.